# Concentration of discrepancy–based approximate Bayesian computation via Rademacher complexity

### v4

@@ 4 Asymptotic properties of ABC with maximum mean discrepancy
As for Proposition 6.1, also Theorem 6.1 shows that informative concentration inequalities similar to those derived in Section 3.2, can be obtained beyond the i.i.d. setting. These results provide comparable insights to those in Theorem 3.2 with the only difference that in this case we require $\sqrt{n}\bar{\varepsilon}_{n}^{2}\to\infty$ rather than $n\bar{\varepsilon}_{n}^{2}\to\infty$ and the term $2b^{2}/n$ within the bound in Theorem 3.2 is now replaced by $2b^{2}/s_{n}$ with $s_{n}\sim\sqrt{n/2}$ as $n\to\infty$. This means that $\bar{\varepsilon}_{n}$ must shrink to zero with a rate at least $n^{1/4}$ slower than the one allowed in the i.i.d. setting. This is an interesting result which clarifies that by moving beyond i.i.d. regimes concentration can still be achieved, although with a slower rate. Such a rate might be pessimistic in some models and we believe it may be improved under future refinements of Lemma 6.1.
Notice that Assumption (VII) could be relaxed to include $\beta$–mixing processes whose coefficients $\beta(k)$ vanish to zero, but at a non–exponential rate, for example, $\beta(k)\sim 1/(k+1)^{\xi}$ for some $\xi>0$. In this case, we could still use Lemma 6.1 to prove concentration, but with a smaller $s_{n}$, which would lead to an even slower concentration. However, we did not provide the most general result for the sake of readability. As for processes that are not $\beta$–mixing, we are not aware of results similar to Lemma 6.1 in this context. This is an important direction for future research.
7 Illustrative simulation in non–i.i.d. settings
We conclude by illustrating the results in Section 6 on a simple simulation study based on a contaminated Gaussian $\textsc{ar}(1)$ process. More specifically, the uncontaminated data are generated from the model $y^{*}_{t}=0.5y^{*}_{t-1}+\varepsilon_{t}$ for $t=1,\ldots,100$ with $\varepsilon_{t}\sim\mbox{N}(0,1)$ independently, and initial state $y^{*}_{0}\sim\mbox{N}(0,1)$. Then, similarly to the simulation study in Section 5, these data are contaminated with a growing fraction $\alpha\in\{0.05,0.10,0.15\}$ of independent realizations from a $\mbox{N}(20,1)$. As such, each observed data point $y_{t}$ is either equal to $y_{t}^{*}$ or to a sample from $\mbox{N}(20,1)$, for $t=1,\ldots,100$. For Bayesian inference, we assume a Gaussian $\textsc{ar}(1)$ model $z_{t}=\theta z_{t-1}+\varepsilon_{t}$ for $t=1,\ldots,100$ with $\varepsilon_{t}\sim\mbox{N}(0,1)$, independently, and focus on learning $\theta$ via discrepancy–based abc under a uniform prior in $[-1,1]$ for such a $\theta$.
Table 2: Concentration and runtimes in seconds (for a single discrepancy evaluation) of abc under mmd with Gaussian kernel, Wasserstein distance, summary–based distance (covariance) and kl divergence for the $\textsc{ar}(1)$ Huber contamination model with a varying $\alpha\in\{0.05,0.10,0.15\}$. $\textsc{mse}=\hat{\mathbb{E}}_{\mu^{*}}[\hat{\mathbb{E}}_{\textsc{abc}}(\theta-\theta_{0})^{2}]$.
@@
Rejection abc is implemented under the same settings and discrepancies considered in Section 5. However, as discussed in Section 6, in this case we focus on distances among the empirical distributions of the $n=m=100$ observed $(y_{0},y_{1}),(y_{1},y_{2}),\ldots,(y_{99},y_{100})$ and synthetic $(z_{0},z_{1}),(z_{1},z_{2}),\ldots,(z_{99},z_{100})$ pairs. This is consistent with the delayed reconstruction strategy in Bernton et al. (2019) and is motivated by the fact that information on $\theta$ is in the bivariate distributions, rather than the marginals. For the same reason, in implementing summary–based abc we consider the sample covariance rather than the sample mean.
@@
Table 2 summarizes the concentration achieved by the different discrepancies analyzed under the aforementioned non–i.i.d. data generating process and model, at varying contamination $\alpha\in\{0.05,0.10,0.15\}$. The results are coherent with those displayed in Table 1 for the i.i.d. scenario and further clarify that discrepancies with guarantees of uniform convergence and concentration generally provide a robust choice, including in non–i.i.d. contexts.
8 Discussion
This article yields important theoretical advancements with respect to the recent literature on the limiting properties of discrepancy–based abc posteriors by connecting these properties with the asymptotic behavior of the Rademacher complexity associated with the chosen discrepancy. As clarified in Sections 2, 3, 4 and 6, although the concept of Rademacher complexity had never been considered within abc before, this notion yields a powerful and promising framework to derive general, informative and uniform limiting properties of discrepancy–based abc posteriors.
While the above contribution already provides key advancements, the proposed perspective based on Rademacher complexity has broader scope and sets the premises for additional future research. For example, as clarified in Sections 24 and 6, any novel result and bound on the Rademacher complexity of specific discrepancies can be directly applied to abc theory through our framework. This may yield tighter or more explicit bounds, possibly holding under milder assumptions and more general discrepancies. For example, to our knowledge, informative bounds for the Rademacher complexity of the Wasserstein distance are currently available only for bounded $\mathcal{Y}$ and, hence, it would be of interest to leverage future findings on the unbounded $\mathcal{Y}$ case to broaden the range of models for which our theory, when specialized to Wasserstein distance, applies. To this end, it might also be promising to explore available results on local Rademacher complexities (e.g., Bartlett et al., 2005), along with the proposed variable–transformation strategy in Example 3.1. Such a latter solution requires the choice of a mapping $g(\cdot)$ which clearly influences the learning properties of the induced distance and, as such, requires further investigation. Finally, although Section 6 provides an important extension beyond the i.i.d. setting, other relaxations of this assumption could be of interest, such as, for instance, the case of independent but not identically distributed data. This setting could be arguably addressed via the residual reconstruction strategy (e.g., Bernton et al., 2019, Section 4.2.3.) that would imply studying discrepancies among empirical distributions of residuals for which the i.i.d. assumption can be again reasonably made.
We shall also emphasize that, although our focus is on the actively–studied convergence and concentration theory for discrepancy–based abc, other properties such as accuracy in uncertainty quantification of credible intervals and the limiting shapes of abc posteriors, in correctly–specified models, have attracted recent interest in the context of summary–based abc (e.g., Frazier et al., 2018), and in a few summary–free implementations (e.g., Frazier, 2020; Wang et al., 2022). While this direction goes beyond the scope of our article, extending and unifying such results, as done for concentration properties, is worth future studies.
Although the ips class is broad, it does not cover all the discrepancies employed in abc. For example, the kl divergence (Jiang et al., 2018) and the Hellinger distance (Frazier, 2020) are not ips but rather belong to the class of $f$–divergences. While this latter family has important differences relative to the ips class, it is worth investigating $f$–divergences in the light of the results in Sections 24 and 6. To accomplish this goal, a possible direction is to exploit the recent unified treatments of these two classes in, e.g., Agrawal and Horel (2021) and Birrell et al. (2022). More generally, our results could also stimulate methodological and theoretical advancements even beyond discrepancy–based abc, especially within generalized likelihood–free Bayesian inference via discrepancy–based pseudo–posteriors (e.g., Bissiri et al., 2016; Jewson et al., 2018; Miller and Dunson, 2019; Chérief-Abdellatif & Alquier, 2020; Matsubara et al., 2022).
Finally, notice that — for the sake of simplicity and ease of comparison with related studies — we have focused on rejection abc, and we have constrained the number $m$ of synthetic samples to be equal to the sample size $n$ of the observed data. While these settings are standard in state–of–the–art theory (Bernton et al., 2019; Frazier, 2020), other abc routines and alternative scenarios where $m$ grows, e.g., sub–linearly, with $n$ deserve further investigation. This latter regime would be especially of interest in settings where the simulation of synthetic data is computationally expensive.
Supplementary Materials
Appendix A Additional Integral Probability Semimetrics
A.1 Two additional remarkable examples of ips discrepancies
While mmd, Wasserstein and summary–based distances provide the most noticeable examples of ips discrepancies employed in abc, two other relevant ips instances are the total variation distance and the Kolmogorov–Smirnov distance, discussed below.
Example A.1.1.
(Total variation distance). Although the total variation (tv) distance is not a common choice within discrepancy–based abc, it still provides a notable example of ips, obtained when $\mathfrak{F}$ is the class of measurable functions whose sup–norm is bounded by $1$; i.e. $\mathfrak{F}=\{f:||f||_{\infty}\leq 1\}$.
Example A.1.2.
A.2 Validity of (III)–(IV) for tv distance and Kolmogorov–Smirnov distance
Examples A.2.1A.2.2 verify in detail the validity of assumptions (III) and (IV) under the tv distance and the Kolmogorov–Smirnov distance, respectively.
Example A.2.1.
(Total variation distance). As mentioned in Example A.1.1 the tv distance is not a common choice in discrepancy–based abc implementations. Such a distance satisfies (III) by definition, but generally not Assumption (IV), unless the cardinality $|\mathcal{Y}|$ of $\mathcal{Y}$ is finite. In fact, when $\mathcal{Y}=\mathbb{R}$ and $\mu\in\mathcal{P}(\mathcal{Y})$ is continuous, the probability that there exists an index $i\neq i^{\prime}$ such that $x_{i}=x_{i^{\prime}}$, is zero. Hence, with probability $1$, for any vector $\epsilon_{1:n}$ of Rademacher variables there always exists a function $f_{\epsilon}$ from $\mathcal{Y}$ to $\{0;1\}$ such that $f_{\epsilon}(x)=\mathds{1}_{\{\epsilon_{i}=1\}}$. Therefore, $\sup_{f\in\mathfrak{F}}|(1/n)\sum_{i=1}^{n}\epsilon_{i}f(x_{i})|\geq(1/n)\sum_{i=1}^{n}\mathds{1}_{\{\epsilon_{i}=1\}}$, which implies that the Rademacher complexity $\mathfrak{R}_{\mu,n}(\mathfrak{F})$ is bounded below by $(1/n)\sum_{i=1}^{n}\mathbb{P}(\epsilon_{i}=1)=1/2.$ Nonetheless, as mentioned above, the tv distance can still satisfy (IV) in specific contexts. For instance, leveraging the bound in Lemma 5.2 of Massart (2000), when the cardinality $|\mathcal{Y}|$ of $\mathcal{Y}$ is finite, there will be replicates in $[f(x_{1}),\ldots,f(x_{n})]$ whenever $n>|\mathcal{Y}|$. Hence, as $n\rightarrow\infty$, it will be impossible to find a function in $\mathfrak{F}$ which can interpolate any noise vector of Rademacher variables with $[f(x_{1}),\ldots,f(x_{n})]$, thus ensuring $\mathfrak{R}_{n}\rightarrow 0$.
Example A.2.2.
(Kolmogorov–Smirnov distance). The ks distance meets (III) by definition and, similarly for mmd with bounded kernels, also condition (IV) is satisfied without the need to impose additional constraints on the model $\mu_{\theta}$ or on the data–generating process. More specifically, Assumption (IV) follows from the inequality $\mathfrak{R}_{\mu,n}(\mathfrak{F})\leq 2[\log(n+1)/n]^{1/2}$ in Chapter 4.3.1 of Wainwright (2019). This is a consequence of the bounds on $\mathfrak{R}_{\mu,n}(\mathfrak{F})$ when $\mathfrak{F}$ is a class of $b$–uniformly bounded functions such that, for some $\nu\geq 1$, it holds $\mbox{card}\{f(x_{1:n}):f\in\mathfrak{F}\}\leq(n+1)^{\nu}$ for any $n$ and $x_{1:n}$ in $\mathcal{Y}^{n}$. When $\mathfrak{F}=\{1_{(-\infty,a]}\}_{a\in\mathbb{R}}$ each $x_{1:n}$ would divide the real line in at most $n+1$ intervals and every indicator function within $\mathfrak{F}$ will take value $1$ for all $x_{i}\leq a$ and zero otherwise, meaning that $\mbox{card}\{f(x_{1:n}):f\in\mathfrak{F}\}\leq(n+1)$. Therefore, by applying equation (4.24) in Wainwright (2019), with $b=1$ and $\nu=1$, yields $\mathfrak{R}_{\mu,n}(\mathfrak{F})\leq 2[\log(n+1)/n]^{1/2}$ for any $\mu\in\mathcal{P}(\mathcal{Y})$, which implies that Assumption (IV) is met. These derivations clarify the usefulness of the available techniques for upper bounding the Rademacher complexity (e.g., Wainwright, 2019, Chapter 4.3), leveraging, in this case, the notion of polynomial discrimination and the closely–related vc dimension.
Appendix B Proofs of Theorems, Corollaries and Propositions
@@
Note that, by leveraging the first inequality in Lemma 2.1, we have $\mathbb{P}_{y_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})>2\mathfrak{R}_{\mu^{*},n}(\mathfrak{F})+\delta]\leq\exp(-n\delta^{2}/2b^{2}).$ Hence, setting $\delta=1/n^{1/4}$, and recalling that $\mathfrak{R}_{\mu^{*},n}(\mathfrak{F})\leq\mathfrak{R}_{n}(\mathfrak{F})$, it follows $\mathbb{P}_{y_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})>2\mathfrak{R}_{n}(\mathfrak{F})+1/n^{1/4}]\leq\exp(-\sqrt{n}/2b^{2})$; note that $\sum_{n\geq 0}\exp(-\sqrt{n}/2b^{2})<\infty$. As a direct consequence of this result, if we define the event
$\mbox{E}_{n}=\{\theta:\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq 2\mathfrak{R}_{n}(\mathfrak{F})+1/n^{1/4}\},$ (B.1)
@@
Hence, in the following we will focus on $\pi_{n}^{(\varepsilon)}\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\}\mathds{1}\{\mbox{E}_{n}\}$. To study such a quantity, recall that, by definition,
@@
Applying again Lemma 2.1 yields $\mathbb{P}_{z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>2\mathfrak{R}_{\mu_{\theta},n}(\mathfrak{F})+\delta\right]\leq\exp(-n\delta^{2}/2b^{2})$. Therefore, setting $\delta=1/n^{1/4}$, and recalling that $\mathfrak{R}_{\mu_{\theta},n}(\mathfrak{F})\leq\mathfrak{R}_{n}(\mathfrak{F})$ and that we are on the event given in (B.1), it follows
$-\exp(-\sqrt{n}/2b^{2})+\mathds{1}\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon-4\mathfrak{R}_{n}(\mathfrak{F})-2/n^{1/4}\}\\ \leq p_{n}(\theta)\leq\exp(-\sqrt{n}/2b^{2})+\mathds{1}\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon+4\mathfrak{R}_{n}(\mathfrak{F})+2/n^{1/4}\}.$ (B.2)
Now, notice that the acceptance probability is defined as $p_{n}=\int p_{n}(\theta)\pi(d\theta)$. Hence, integrating with respect to $\pi(\theta)$ in the above inequalities yields, for $n\rightarrow\infty$,
$\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon-c_{\mathfrak{F}}\}+o(1)\leq p_{n}\leq\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon+c_{\mathfrak{F}}\}+o(1),$
where $c_{\mathfrak{F}}=4\lim\sup\mathfrak{R}_{n}(\mathfrak{F})$, as in equation (3.1), thus concluding the first part of the proof.
To proceed with the second part of the proof, notice that, by the definition of $\tilde{\varepsilon}=\inf\{\varepsilon>0:\smash{\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\tilde{\varepsilon}\}}>0\}$, the left part of the above inequality is bounded away from zero for $n$ large enough, whenever $\varepsilon-c_{\mathfrak{F}}>\tilde{\varepsilon}$. This implies that also the acceptance probability $p_{n}$ is strictly positive. As a consequence, for such $n$, it follows that
@@
is well–defined for any event A. Then, leveraging the upper bound in (B.2) yields
@@
Note that by combining Equation (2.2) in Lemma 2.1 with the result $\sum_{n>0}\exp[-n\delta^{2}/(2b^{2})]<\infty$, the Borel–Cantelli Lemma implies that both $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ and $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ converge to 0 almost surely when $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$ as $n\rightarrow\infty$. Hence, since
@@
Combining the above result with the proof of Theorem 1 in Jiang et al. (2018) yields the statement of Corollary 3.1. Notice that, as discussed in Section 3.1, the limiting pseudo–posterior in Corollary 3.1 is well–defined only for those $\varepsilon\geq\tilde{\varepsilon}$, with $\tilde{\varepsilon}$ as in Theorem 3.1. ∎
@@
Since Lemma 2.1 and $\mathfrak{R}_{n}(\mathfrak{F})=\sup_{\mu\in\mathcal{P}(\mathcal{Y})}\mathfrak{R}_{\mu,n}(\mathfrak{F})\geq\mathfrak{R}_{\mu,n}(\mathfrak{F})$ hold for every $\mu\in\mathcal{P}(\mathcal{Y})$, then, for every integer $n\geq 1$ and any scalar $\delta\geq 0$, Equation (2.2) implies $\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{n}(\mathfrak{F})+\delta\right]\geq 1-\exp(-n\delta^{2}/2b^{2}).$ Moreover, since this result holds for any $\delta\geq 0$, it follows that $\mathbb{P}_{x_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{n}(\mathfrak{F})+(c_{1}-2\mathfrak{R}_{n}(\mathfrak{F}))]\geq 1-\exp[-n(c_{1}-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}],$ for any $c_{1}\geq 2\mathfrak{R}_{n}(\mathfrak{F})$. Hence,
$\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq c_{1}\right]\geq 1-\exp[-n(c_{1}-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}].$ (B.3)
Recalling the settings of Theorem 3.2, consider the sequence $\bar{\varepsilon}_{n}\to 0$ as $n\to\infty$, with $n\bar{\varepsilon}^{2}_{n}\to\infty$ and $\bar{\varepsilon}_{n}/\mathfrak{R}_{n}(\mathfrak{F})\to\infty$, which is possible by the Assumption (IV). These regimes imply that $\bar{\varepsilon}_{n}$ goes to zero slower than $\mathfrak{R}_{n}(\mathfrak{F})$ and, therefore, for $n$ large enough, $\bar{\varepsilon}_{n}/3>2\mathfrak{R}_{n}(\mathfrak{F})$. Therefore, under Assumptions (I)–(III) it is now possible to apply (B.3) to $y_{1:n}$, by setting $c_{1}=\bar{\varepsilon}_{n}/3$, which yields

### v5

@@ Appendix B Concentration in the space of parameters
As for Proposition C.6, also Theorem C.7 shows that informative concentration inequalities similar to those derived in Section 3.2, can be obtained beyond the i.i.d. setting. These results provide insights comparable to those in Theorem 3.3 with the only difference that in this case we require $\sqrt{n}\bar{\varepsilon}_{n}^{2}\to\infty$ rather than $n\bar{\varepsilon}_{n}^{2}\to\infty$ and the term $2b^{2}/n$ within the bound in Theorem 3.3 is now replaced by $2b^{2}/s_{n}$ with $s_{n}\sim\sqrt{n}/2$ as $n\to\infty$. This means that $\bar{\varepsilon}_{n}$ must shrink to zero with a rate at least $n^{1/4}$ slower than the one allowed in the i.i.d. setting. This is an interesting result which clarifies that when moving beyond i.i.d. regimes concentration can still be achieved, although with a slower rate. Such a rate might be pessimistic in some models and we believe it may be improved under future refinements of Lemma C.2.
Notice that (VII) could be relaxed to include $\beta$-mixing processes whose coefficients $\beta(k)$ vanish to zero, but at a non-exponential rate, e.g., $\beta(k)\sim 1/(k+1)^{\xi}$ for some $\xi>0$. In this case, we could still use Lemma C.2 to prove concentration, but with a smaller $s_{n}$, that would lead to even slower rates. However, we did not provide the most general result for the sake of readability. As for processes that are not $\beta$-mixing, we are not aware of results similar to Lemma C.2 in this context. This is an important direction for future research.
C.2 Illustrative simulation in non-i.i.d. settings
Let us illustrate the results in Section C.1 on a simple simulation study focusing on a contaminated Gaussian $\textsc{ar}(1)$ process. More specifically, the uncontaminated data are generated from the model $y^{*}_{t}=0.5y^{*}_{t-1}+\varepsilon_{t}$ for $t=1,\ldots,100$ with $\varepsilon_{t}\sim\mbox{N}(0,1)$ independently, and initial state $y^{*}_{0}\sim\mbox{N}(0,1)$. Then, similarly to the simulation study in Section 5, these data are contaminated with a growing fraction $\alpha\in\{0.05,0.10,0.15\}$ of independent realizations from a $\mbox{N}(20,1)$. As such, each observed data point $y_{t}$ is either equal to $y_{t}^{*}$ or to a sample from $\mbox{N}(20,1)$, for $t=1,\ldots,100$. For Bayesian inference, we assume an $\textsc{ar}(1)$ model $z_{t}=\theta z_{t-1}+\varepsilon_{t}$, with $\varepsilon_{t}\sim\mbox{N}(0,1)$, and focus on learning $\theta$ via discrepancy-based abc under a uniform prior on $[-1,1]$ for $\theta$.
Table C.1: Concentration and runtimes in seconds (for a single discrepancy evaluation) of abc under mmd with Gaussian kernel, Wasserstein-1 distance, summary-based distance (covariance) and kl divergence for an $\textsc{ar}(1)$ Huber contamination model with $\alpha\in\{0.05,0.10,0.15\}$. $\textsc{mse}=\hat{\mathbb{E}}_{\mu^{*(n)}}[\hat{\mathbb{E}}_{\textsc{abc}}(\theta-\theta_{0})^{2}]$, $\theta_{0}=0.5$.
@@
Rejection abc is implemented under the same settings and discrepancies considered in Section 5. However, as discussed in Section C.1, in this case we focus on distances among the empirical distributions of the $n=m=100$ observed $(y_{0},y_{1}),(y_{1},y_{2}),\ldots,(y_{99},y_{100})$ and synthetic $(z_{0},z_{1}),(z_{1},z_{2}),\ldots,(z_{99},z_{100})$ pairs. This is consistent with the delay reconstruction strategy in Bernton et al. (2019) and is motivated by the fact that information on $\theta$ is in the bivariate distributions, rather than in the marginals. For the same reason, in implementing summary-based abc we consider the sample covariance rather than the sample mean.
@@
Table C.1 summarizes the concentration achieved by the different discrepancies analyzed under the aforementioned non-i.i.d. data generating process and model, at varying contamination $\alpha\in\{0.05,0.10,0.15\}$. The results are coherent with those displayed in Table 1 for the i.i.d. scenario and further clarify that discrepancies with guarantees of uniform convergence and concentration generally provide a robust choice, including in non-i.i.d. contexts.
Appendix D Proofs of Theorems, Corollaries and Propositions
@@
Note that, by leveraging the first inequality in Lemma 2.6, we have $\mathbb{P}_{y_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})>2\mathfrak{R}_{\mu^{*},n}(\mathfrak{F})+\delta]\leq\exp(-n\delta^{2}/2b^{2}).$ Hence, setting $\delta=1/n^{1/4}$, and recalling that $\mathfrak{R}_{\mu^{*},n}(\mathfrak{F})\leq\mathfrak{R}_{n}(\mathfrak{F})$, it follows $\mathbb{P}_{y_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})>2\mathfrak{R}_{n}(\mathfrak{F})+1/n^{1/4}]\leq\exp(-\sqrt{n}/2b^{2})$; note that $\sum_{n\geq 0}\exp(-\sqrt{n}/2b^{2})<\infty$. Therefore, if we define the event
$\mbox{E}_{n}=\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq 2\mathfrak{R}_{n}(\mathfrak{F})+1/n^{1/4}\},$ (D.1)
@@
Hence, in the following we focus on $\pi_{n}^{(\varepsilon)}\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\}\mathds{1}\{\mbox{E}_{n}\}$. To this end, recall that
@@
Applying again Lemma 2.6 yields $\mathbb{P}_{z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>2\mathfrak{R}_{\mu_{\theta},n}(\mathfrak{F})+\delta\right]\leq\exp(-n\delta^{2}/2b^{2})$. Therefore, setting $\delta=1/n^{1/4}$, and recalling that $\mathfrak{R}_{\mu_{\theta},n}(\mathfrak{F})\leq\mathfrak{R}_{n}(\mathfrak{F})$ and that we are on the event given in (D.1), it follows
$-\exp(-\sqrt{n}/2b^{2})+\mathds{1}\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon-4\mathfrak{R}_{n}(\mathfrak{F})-2/n^{1/4}\}\\ \leq p_{n}(\theta)\leq\exp(-\sqrt{n}/2b^{2})+\mathds{1}\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon+4\mathfrak{R}_{n}(\mathfrak{F})+2/n^{1/4}\}.$ (D.2)
Now, notice that the acceptance probability is defined as $p_{n}=\int p_{n}(\theta)\pi(d\theta)$. Hence, integrating with respect to $\pi(\theta)$ in the above inequalities yields, for $n$ large enough,
$\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon-c_{\mathfrak{F}}\}-e_{n}\leq p_{n}\leq\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon+c_{\mathfrak{F}}\}+e_{n},$
where $c_{\mathfrak{F}}=4\lim\sup\mathfrak{R}_{n}(\mathfrak{F})$, as in Equation (3), thus concluding the first part of the proof.
To proceed with the second part of the proof, notice that, by the definition of $\tilde{\varepsilon}=\inf\{\epsilon>0:\smash{\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\epsilon\}}>0\}$, the left part of the above inequality is bounded away from zero for $n$ large enough, whenever $\varepsilon-c_{\mathfrak{F}}>\tilde{\varepsilon}$. This implies that also the acceptance probability $p_{n}$ is strictly positive. As a consequence, for such $n$, it follows that
@@
is well-defined for any event A. Then, leveraging the upper bound in (D.2) yields
@@
Note that by combining Equation (2) in Lemma 2.6 with the result $\sum_{n>0}\exp[-n\delta^{2}/(2b^{2})]<\infty$, the Borel–Cantelli Lemma implies that both $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ and $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ converge to 0 almost surely when $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$ as $n\rightarrow\infty$. Hence, since
@@
Combining this result with the proof of Theorem 1 in Jiang, Wu and Wong (2018) yields the statement of Corollary 3.2. Notice that, as discussed in Section 3.1, the limiting pseudo-posterior in Corollary 3.2 is well-defined only for those $\varepsilon>\tilde{\varepsilon}$, with $\tilde{\varepsilon}$ as in Theorem 3.1. ∎
@@
Since Lemma 2.6 and $\mathfrak{R}_{n}(\mathfrak{F})=\sup_{\mu\in\mathcal{P}(\mathcal{Y})}\mathfrak{R}_{\mu,n}(\mathfrak{F})\geq\mathfrak{R}_{\mu,n}(\mathfrak{F})$ hold for every $\mu\in\mathcal{P}(\mathcal{Y})$, then, for every integer $n\geq 1$ and any scalar $\delta\geq 0$, Equation (2) implies $\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{n}(\mathfrak{F})+\delta\right]\geq 1-\exp(-n\delta^{2}/2b^{2}).$ Moreover, since this result holds for any $\delta\geq 0$, it follows that $\mathbb{P}_{x_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{n}(\mathfrak{F})+(c_{1}-2\mathfrak{R}_{n}(\mathfrak{F}))]\geq 1-\exp[-n(c_{1}-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}],$ for any $c_{1}\geq 2\mathfrak{R}_{n}(\mathfrak{F})$. Hence,
$\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq c_{1}\right]\geq 1-\exp[-n(c_{1}-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}].$ (D.3)
Recalling the settings of Theorem 3.3, consider the sequence $\bar{\varepsilon}_{n}\to 0$ as $n\to\infty$, with $n\bar{\varepsilon}^{2}_{n}\to\infty$ and $\bar{\varepsilon}_{n}/\mathfrak{R}_{n}(\mathfrak{F})\to\infty$, which is possible by Assumption (IV). These regimes imply that $\bar{\varepsilon}_{n}$ goes to zero slower than $\mathfrak{R}_{n}(\mathfrak{F})$ and, hence, for $n$ large enough, $\bar{\varepsilon}_{n}/3>2\mathfrak{R}_{n}(\mathfrak{F})$. Therefore, under Assumptions (I)–(III) it is now possible to apply (D.3) to $y_{1:n}$, by setting $c_{1}=\bar{\varepsilon}_{n}/3$, which yields